OpenStudy (smurfy14):

|√(5)-(5/2)| (abs value of sq rt of 5 minus the fraction 5 over 2)

OpenStudy (anonymous):

This?\[\left|\sqrt{5} - \frac{5}{2}\right|\]

OpenStudy (smurfy14):

yes. can you show me how to do it?

OpenStudy (anonymous):

What are you trying to do with it?

OpenStudy (smurfy14):

rewrite the expression without using abs value signs.

OpenStudy (anonymous):

Oh. Well which one is greater?

OpenStudy (anonymous):

\(\sqrt{5}\) or \( \large 5\over 2\)

OpenStudy (smurfy14):

isnt it -5/2? if not then 5/2

OpenStudy (anonymous):

Ok so\[\sqrt{5} < \frac{5}{2} \implies \sqrt{5} - \frac{5}{2} < 0\] So if the part inside the absolute value is negative what do we do?

OpenStudy (smurfy14):

change it to positive

OpenStudy (anonymous):

And how do you make a negative number into a positive number?

OpenStudy (smurfy14):

idk :(

OpenStudy (smurfy14):

oh divide by -1?

OpenStudy (anonymous):

Sure, or multiply by -1. So \[\left|\sqrt{5} - \frac{5}{2}\right| = -\left(\sqrt{5} - \frac{5}{2}\right)\] Then simplify that.

OpenStudy (smurfy14):

so the answer is sq rt 5+5/2?